OpenStudy (anonymous):

I am trying to understand example 1.7 on "Differential Equations Desmistified" book but I can't understand how the author convert (x/2)*cot(y)dy/dx=-1 into 2X*sin(y)dx+x^2*cos(y)dy=0..Can you please explain how? thank you very much

OpenStudy (mathstudent55):

cot(y) = cos(y)/sin(y)

OpenStudy (dumbcow):

\[\frac{x \cos(y) dy}{2 \sin(y) dx} + 1 = 0\] combine into single fraction \[\frac{x \cos(y) dy + 2 \sin(y) dx}{2 \sin(y) dx} = 0\] ...

OpenStudy (dumbcow):

oh and then multiply equation by "x"

OpenStudy (mathstudent55):

\(\dfrac{x}{2} \cot y \dfrac{dy}{dx}=-1\) \(\dfrac{x}{2} \dfrac{\cos y}{\sin y} \dfrac{dy}{dx}=-1\) Multiply both sides by \(2 \sin y ~dx\) \(x \cos y dy = -2 \sin y ~dx\) Add \(2 \sin y ~dx\) to both sides. \(2 \sin y ~dx + x \cos y ~dy = 0\) Multiply both sides by x: \(2x \sin y ~dx + x^2 \cos y ~dy = 0\)

OpenStudy (anonymous):

Thank you very much!!! to all of you! I just discovered this site and I love it.

OpenStudy (mathstudent55):

You are welcome.